On the parallel line (y-3=3(x+1)) find the ordered pair where x=2
Answer by Alan3354(69443)

On the parallel line (y-3=3(x+1)) find the ordered pair where x=2
--------------
What parallel line?
---
(y-3=3(x+1)) find the ordered pair where x=2
-----
(y-3=3(x+1))
Sub 2 for x.
y-3=3(2+1) = 9
y = 12
--> (2,12)